OpenStudy (anonymous):

Solve the equation: 3/(x+1) - 1/x = 1/2

OpenStudy (anonymous):

\[\frac{3}{x+1}-\frac{1}{x}=\frac{1}{2}\]

OpenStudy (anonymous):

best is to multiply all by \(2x(x+1)\) to clear the fractions.

OpenStudy (anonymous):

whoa mistake!!

OpenStudy (anonymous):

you get \[3(2x)-2(x+1)=x(x+1)\] i forgot to cancel the 2

OpenStudy (anonymous):

ok from here? you get a quadratic equation that is easy to factor

OpenStudy (anonymous):

Just multiply both sides of the equation with 2x(x+1). Pretty easy.